The area of a rectangular floor is 105 square feet. if its length is 1 more than twice its width, find the length and width.
Found 2 solutions by fractalier, Boreal:
Answer by fractalier(6550)

Area A = LW = 105
We also have L = 2W + 1.
Now substitute and get
(2W+1)W = 105
2W^2 + W - 105 = 0
Now factor and solve
(2W + 15)(W - 7) = 0
so that
W = 7 feet and
L = 2W + 1 = 15 feet

Answer by Boreal(15235)

width=x
length=2x+1
area = (2x+1)(x)=2x^2+x=105
2x^2+x-105=0
(2x+15)(x-7)=0
x=7, -7.5, use positive root only
2x+1=15
it is 7 feet x 15 feet